how many cubes 6in on an edge will fit into a bin that is 24in on an edge?

Answer by checkley79(3341)

EACH 6IN. CUBE IS 6^3=216 IN^3
THE BIN IS 24^3=13,824 IN^2
13,824/216=64 6 IN. CUBES WILL FIT IN THE BIN